# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

### v2

@@ 5.1 Asymptotic Normality
We next provide sufficient conditions for consistency and asymptotic normality of the estimator $\hat{\psi}^{mr}$.
@@
    There exists some constant $C$ such that $\|\mu_{1}^{2}(M,X)\|_{2}$, $\|\mu_{2}^{2}(X)\|_{2}$, and $\mathbb{E}[Y^{4}]$, are upper bounded by $C$.
@@
For any choices of the nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, define,
@@
We have the following result regarding the asymptotic behavior of $\hat{\psi}^{mr}$.
@@
Under Assumptions 2 and  3, the estimator $\hat{\psi}^{mr}$ is $\sqrt{n}$-consistent and asymptotically normal and satisfies
$\sqrt{n}(\hat{\psi}^{mr}-\psi_{0})=\sqrt{n}\mathbb{E}_{n}\left[\Phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2},\psi_{0})\right]+o_{p}(1).$
@@
$\hat{\sigma}_{\ell}^{2}=\frac{1}{m}\sum_{i=1}^{m}\Phi^{2}(O_{i,\ell};\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell},\hat{\psi}_{\ell}^{mr}),$
where $O_{i,\ell}\in I_{\ell}$, and define,
@@
$\hat{\psi}^{mr}\pm z_{1-\alpha/2}\frac{\hat{\sigma}}{\sqrt{n}},$
@@ 6 Conclusion
In estimating a functional of interest that involves nuisance functions, if one naively aims to perfectly fit the nuisance functions, the resulting estimator of the parameter of interest can be sensitive to even small errors in the nuisance function estimators. Therefore, in order to estimate a functional, it is important to investigate the role of nuisance functions involved in that functional in the structure of the bias. Based on this vision, we proposed a new approach for estimating the nuisance functions of the mediation functional when the influence function-based estimator is used. Our proposed estimators are based on the approach of balancing that can target the bias of the parameter of interest more directly and also can lead to an estimator which is less sensitive to error. We demonstrated that this approach retains the multiple-robustness property. Moreover, we provided conditions on our nuisance function estimators sufficient for consistency and asymptotic normality of the influence function-based estimator of the mediation functional.
@@ Appendix A Proofs
$\arg\max_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}^{2}+\frac{1}{4}h^{2}_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]$
$=\arg\min_{h_{1}}\mathbb{E}\left[\left(\frac{1}{2}h_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}.$
@@
$\breve{\pi}_{1}$ $=\arg\min_{\pi_{1}}\max_{h_{1}}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[2\left\{(1-A)\pi_{1}(X)-1\right\}\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}-\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}^{2}\right]$
$=\arg\min_{\pi_{1}}\mathbb{E}\left[\mathbb{E}\bigg{[}(1-A)\pi_{1}(X)-1\bigg{|}X\bigg{]}^{2}\right],$
@@
$\mathbb{E}\bigg{[}(1-A)\breve{\pi}_{1}(X)-1\bigg{|}X\bigg{]}=0.$
@@
$\mathbb{E}\left[\left\{(1-A)\breve{\pi}_{1}(X)-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\mathbb{E}\bigg{[}(1-A)\breve{\pi}_{1}(X)-1\bigg{|}X\bigg{]}h_{1}(X)\right]=0.$
@@
$\arg\max_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)(1-A)\hat{\pi}_{1}(X)\right\}^{2}+\frac{1}{4}h^{2}_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)\right]$
$=\arg\min_{h_{2}}\mathbb{E}\left[\left(\frac{1}{2}h_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\right)^{2}\right]$
$=2\mathbb{E}\bigg{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}.$
@@
$\breve{\pi}_{2}$ $=\arg\min_{\pi_{2}}\max_{h_{2}}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{\pi_{2}}\mathbb{E}\bigg{[}2\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\mathbb{E}\big{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\big{|}M,X\big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathbb{E}\big{[}A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\big{|}M,X\big{]}^{2}\bigg{]}$
$=\arg\min_{\pi_{2}}\mathbb{E}\left[\mathbb{E}\bigg{[}A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}^{2}\right],$
@@
$\mathbb{E}\bigg{[}A\breve{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg{|}M,X\bigg{]}=0.$
@@
$\mathbb{E}\left[\left\{A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
$=\mathbb{E}\left[\mathbb{E}\bigg{[}A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\bigg{|}M,X\bigg{]}h_{2}(M,X)\right]=0.$
@@
$\max_{h_{1}\in\mathcal{H}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
=$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
@@
$\mathbb{E}_{n}[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)]$ $=\sum_{i=1}^{n}h(X_{i})\{\xi_{n}(\pi_{1})\}_{i}={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\},$
$\mathbb{E}_{n}[h_{1}^{2}(X)]$ $=\frac{1}{n}{\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}^{2}{\alpha^{\ast}},$
@@
$\mathbb{E}_{n}\Big{[}\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big{]}-\frac{1}{4}\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\}-{\alpha^{\ast}}^{\top}(\frac{1}{4n}K_{\mathcal{H}_{1},n}^{2}+\frac{1}{4}\lambda_{\mathcal{H}_{1}}K_{\mathcal{H}_{1},n}){\alpha^{\ast}}.$
@@
$\alpha^{\ast}=2(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
@@
$\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\Big{[}\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big{]}-\frac{1}{4}\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
=$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\}.$
@@
where $\Gamma_{1}=K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$
@@
$=\min_{\alpha\in\mathbb{R}^{n}}\frac{1}{n^{2}}\alpha^{\top}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\frac{2}{n^{2}}e_{n}^{\top}\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\lambda_{\Pi_{1}}\alpha^{\top}K_{\Pi_{1},n}\alpha+c,$
@@
$\alpha=-\Big{(}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big{)}^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n}.$
@@
    $\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=\mathbb{E}\left[\left\{\mathbb{E}[1-A\mid X]\pi_{1}(X)-1\right\}h_{1}(X)\right]=0.$

### v3

@@ 5.2 Asymptotic Normality
We next provide sufficient conditions for root-$n$ consistency and asymptotic normality of the estimator $\hat{\psi}^{\text{2S}}$.
@@
    There exists some constant $C$ such that $\|\mu_{1}^{2}(M,X)\|_{2}$, $\|\mu_{2}^{2}(X)\|_{2}$, and $E[Y^{4}]$, are upper bounded by $C$.
@@
For any choice of the nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, define,
@@
We have the following result regarding the asymptotic behavior of $\hat{\psi}^{\text{2S}}$.
@@
Under Assumptions 2 and  3, the estimator $\hat{\psi}^{\text{2S}}$ is $\sqrt{n}$-consistent and asymptotically normal and satisfies
$\sqrt{n}(\hat{\psi}^{\text{2S}}-\psi_{0})=\sqrt{n}E_{n}\left[\Phi(O;\pi_{1},\pi_{2},\mu_{1},\mu_{2},\psi_{0})\right]+o_{p}(1).$
@@
$\hat{\sigma}_{\ell}^{2}=E_{m}\big[\Phi^{2}(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell},\hat{\psi}^{\text{2S}})\big],$
and define,
@@
$\hat{\psi}^{\text{2S}}\pm z_{1-\alpha/2}\frac{\hat{\sigma}}{\sqrt{n}},$
@@
6 Simulation Study
In this section, we evaluate our proposed method on synthetic data for estimating the mediation functional $\psi_{0}$. We consider a data generating process (DGP) which satisfies Assumptions 1 and 2, and study the robustness property of our estimator under the scenarios specified in Assumption 3. In particular, consider a three-dimensional covariate vector $X=(X_{1},X_{2},X_{3})$. The continuous covariates $X_{1}$ and $X_{2}$ are independently generated from the standard normal distribution $\mathcal{N}(0,1)$, and the binary covariate $X_{3}$ is generated from a Bernoulli distribution with $p(X_{3}=1)=0.5$. Treatment $A$ is generated according to
$p(A=1\mid X_{1},X_{2},X_{3})=\text{expit}(0.35+0.1X_{1}-0.2X_{2}+0.3X_{3}-0.15X_{1}^{2}+0.25X_{2}X_{3}).$
The mediator $M$ is drawn from a two-component Gaussian mixture centered around
$\mu_{M}(X_{1},X_{2},X_{3},A)=2.4+1.6X_{1}-1.2X_{2}+2.2X_{3}+0.8A+1.5X_{1}^{2}-1.8X_{2}X_{3}.$
Specifically,
$M\sim\begin{cases}\mathcal{N}(\mu_{M}(X_{1},X_{2},X_{3},A)-\delta,\sigma^{2}),&\text{with probability }p_{\text{left}},\\ \mathcal{N}(\mu_{M}(X_{1},X_{2},X_{3},A)+\delta,\sigma^{2}),&\text{with probability }1-p_{\text{left}},\end{cases}$
where $\sigma^{2}=1$ is the within-mode variance, $\delta=4$ is half the separation between the two modes, and $p_{\text{left}}=0.5$ controls the mixing proportion. Finally, the outcome $Y$ is generated from a Gaussian distribution with mean function
$\mu_{Y}(X_{1},X_{2},X_{3},M,A)$
$=6.4+2.7X_{1}+0.7X_{2}-3.6X_{3}-0.8X_{2}^{2}+2.5X_{1}A+3.2MA-2.9M+4.5A,$
and variance fixed at 1.
We consider sample sizes $n\in\{1000,2000,4000\}$ and repeat simulations under each sample size for 300 times. Cross-fitting is implemented with four folds, and hyperparameters for estimating $\pi_{1}$ and $\pi_{2}$ are tuned via cross-validation using a 75/25 train-validation split. For comparison, we also present the results for the TTS estimator $\hat{\psi}^{\text{TTS}}$ and a naive estimator $\hat{\psi}^{\text{naive}}=\frac{1}{L}\sum_{\ell=1}^{L}E_{m}[\hat{\mu}_{2,\ell}(X)]$, which is constructed from the identification formula $E\{E[E[Y\mid A=1,M,X]\mid A=0,X]\}$. In our implementation, $\mu_{1}$ and $\mu_{2}$ are estimated via random forests. The naive estimator uses the same $\mu_{2}$, while the TTS estimator also relies on random forests for nuisance estimation, except for the conditional density of the mediator, which is obtained using kernel density estimator. Replication code for the simulation study is available at https://github.com/changliues/TwoStage.git.
Table 1: Results for comparing TTS, naive, and two-stage estimators.
Sample Size Estimators Mean Bias Median Bias RMSE SD
n=1000 $\hat{\psi}^{\text{TTS}}$ $2.17\times 10^{33}$ $0.662\times 10^{-1}$ $3.70\times 10^{34}$ $3.70\times 10^{34}$
$\hat{\psi}^{\text{naive}}$ $1.59\times 10^{-1}$ $1.70\times 10^{-1}$ $2.69\times 10^{-1}$ $2.17\times 10^{-1}$
$\hat{\psi}^{\text{2S}}$ $1.68\times 10^{-1}$ $1.80\times 10^{-1}$ $2.76\times 10^{-1}$ $2.18\times 10^{-1}$
n=2000 $\hat{\psi}^{\text{TTS}}$ $4.15\times 10^{39}$ $2.47\times 10^{-1}$ $7.19\times 10^{40}$ $7.19\times 10^{40}$
$\hat{\psi}^{\text{naive}}$ $1.29\times 10^{-1}$ $1.40\times 10^{-1}$ $1.81\times 10^{-1}$ $1.27\times 10^{-1}$
$\hat{\psi}^{\text{2S}}$ $1.18\times 10^{-1}$ $1.22\times 10^{-1}$ $1.76\times 10^{-1}$ $1.31\times 10^{-1}$
n=4000 $\hat{\psi}^{\text{TTS}}$ $3.70\times 10^{60}$ $7.70\times 10^{-1}$ $1.07\times 10^{149}$ $1.07\times 10^{149}$
$\hat{\psi}^{\text{naive}}$ $13\times 10^{-1} 1.1$$16\times 10^{-1}$ $1.48\times 10^{-1}$ $0.965\times 10^{-1}$
$\hat{\psi}^{\text{2S}}$ $0.686\times 10^{-1}$ $0.696\times 10^{-1}$ $1.20\times 10^{-1}$ $0.986\times 10^{-1}$
Table 2: 95% confidence intervals and coverage for the two-stage estimator.
Sample Size 95% CI Coverage
n=1000 [9.590, 10.346] 0.870
n=2000 [9.643, 10.191] 0.883
n=4000 [9.670, 10.066] 0.910
Table 3: Robustness results for the two-stage estimator: Bias (RMSE).
Sample size all true $\{\pi_{1},\pi_{2}\}$ $\{\mu_{1},\mu_{2}\}$ $\{\pi_{1},\mu_{1}\}$ all false
n=1000 0.168 (0.276) 0.156 (0.276) 0.189 (0.287) 0.163 (0.273) 0.163 (0.273)
n=2000 0.118 (0.176) 0.137 (0.193) 0.184 (0.224) 0.155 (0.202) 0.185 (0.224)
n=4000 0.0686 (0.120) 0.108 (0.149) 0.160 (0.187) 0.140 (0.172) 0.203 (0.226)
Table 1 compares the TTS estimator, naive estimator, and two-stage estimator. The TTS estimator is highly unstable when the mediator is continuous, with mean bias and RMSE becoming excessively large due to the density in the denominator. In contrast, the naive estimator and the two-stage estimator remain stable. However, the proposed two-stage estimator exhibits lower bias and consistently outperforms the naive estimator in large samples. Results in Table 1 demonstrate the consistency of the two stage estimator. Moreover, as shown in Table 2, coverage for the two-stage estimator approaches the nominal 95% as sample size increases.
Table 3 illustrates the robustness property of the two-stage estimator. Each scenario corresponds to a setting in which specific subsets of nuisance functions are estimated well by the random forests, while the others are fit with suboptimal hyperparameters, matching the cases described in Proposition 6. Across the first four scenarios, the two-stage estimator remains stable, with bias decreasing as sample size grows, and it performs best when all nuisance functions are well-estimated. Even in the all false case, the estimator still remains stable. Together with Table 1, this confirms that the proposed two-stage estimator is both more accurate than the other two approaches and robust to nuisance function misspecification.
7 Application: Effect of Obesity on Cardiovascular Disease
In this section, we apply our proposed method to study the direct and indirect effect of obesity on coronary heart disease (CHD) mediated through the Glycohemoglobin (HbA1c). HbA1c, measured as a percentage of long-term blood glucose control, is the standard biomarker for diagnosing and monitoring diabetes, which is itself a major risk factor for CHD (Xu et al.,, 2017). Disentangling the extent to which obesity contributes to CHD risk through HbA1c, as opposed to other biological pathways, is of both scientific and policy relevance. If a substantial portion of the effect is mediated through HbA1c, interventions focused on glycemic control may effectively reduce obesity-related CHD risk; if not, alternative mechanisms such as inflammation, lipid metabolism, or hypertension may be more important to target. Partitioning the total effect into direct and indirect components thus provides critical insight for policymakers in determining whether diabetes prevention and management strategies are sufficient to mitigate CHD burden, or whether obesity itself must be targeted more aggressively regardless of glycemic control.
We use data from the 2013-2014 cycle of National Health and Nutrition Examination Survey (NHANES, National Center for Health Statistics, (2014)) for analysis. Obesity is defined using body mass index (BMI): individuals with BMI $\leq 29$ are classified as non-exposed (non-obese), and those with BMI $\geq 31$ are classified as exposed (obese). To sharpen the contrast between exposure groups, we exclude individuals with BMI in the range 29-31, leaving a deliberate gap around the threshold. The outcome of interest is a binary indicator based on self-reported physician diagnosis, where $Y=1$ indicates that the individual has previously been diagnosed with CHD. We adjust for eleven covariates that are potential confounders: age, gender, race, education level, annual family income, alcohol use, smoking history, vigorous work activity, diet quality, sleeping habits, and health insurance coverage. We perform complete case analysis, yielding a final sample size of 2685 and report effects in terms of odds ratios (OR) for ease of interpretation.
To motivate the definitions of direct and indirect ORs, recall from Section 2 that the ATE can be decomposed into the sum of the NIE and NDE. In an analogous manner, when considering binary outcomes, we define odds ratios in terms of the corresponding counterfactual event odds. The natural direct effect odds ratio is defined as
$\text{OR}_{\text{NDE}}=\frac{p(Y^{(1,M^{(0)})}=1)/(1-p(Y^{(1,M^{(0)})}=1))}{p(Y^{(0,M^{(0)})}=1)/(1-p(Y^{(0,M^{(0)})}=1))},$
the natural indirect effect odds ratio is defined as
$\text{OR}_{\text{NIE}}=\frac{p(Y^{(1,M^{(1)})}=1)/(1-p(Y^{(1,M^{(1)})}=1))}{p(Y^{(1,M^{(0)})}=1)/(1-p(Y^{(1,M^{(0)})}=1))},$
and the total effect odds ratio equals their product, $\text{OR}_{\text{TOT}}=\text{OR}_{\text{NIE}}\times\text{OR}_{\text{NDE}}$. While our earlier parameter of interest, $\theta_{0}=E[Y^{(1,M^{(0)})}]$, was expressed as a mean potential outcome, note that because the outcome is binary, this mean coincides with the probability $p(Y^{(1,M^{(0)})}=1)$, which then serves as the basis for constructing odds ratios. Since the denominator of $\text{OR}_{\text{NDE}}$ and numerator of $\text{OR}_{\text{NIE}}$ do not involve the mediator mechanism, we use the estimator $\frac{1}{L}\sum_{\ell=1}^{L}E_{m}[\hat{p}_{\ell}(Y=1\mid A=a,X)]$ for $a\in\{0,1\}$.
[…]
Taken together, our findings suggest that obesity results in approximately 60% higher odds of CHD compared to non-obesity, and that, considering only the direct pathway independent of HbA1c, obesity leads to about 35% higher odds of CHD.
@@ 8 Conclusion
When estimating a functional of interest that depends on nuisance functions, naively aiming for perfect fits for these nuisance functions can make the resulting estimator highly sensitive, as even small errors in the nuisance function estimators may be amplified, especially in structures involving ratios of density functions or scores in the denominator. Instead, it is important to investigate the role of nuisance functions in the bias structure of the estimator. Based on this vision, we proposed a new approach for estimating the nuisance functions of the mediation functional, tailored to the bias structure of the influence function-based estimator. Our proposed nuisance estimators are based on the approach of balancing that can target the bias of the parameter of interest directly and also can lead to an estimator which is less sensitive to error. We demonstrated that the resulting estimator of the mediation function is multiply robust to the misspecification of the nuisance functions. Moreover, we provided conditions on the nuisance function estimators sufficient for obtaining root-$n$ consistency and asymptotic normality for the estimator of the mediation functional. We evaluated our proposed approach on synthetic data and applied our method to estimate the effect of obesity on coronary heart disease mediated by Glycohemoglobin, finding that obesity results in a 60% higher odds of coronary heart disease, of which roughly 35% is attributable to the direct pathway independent of Glycohemoglobin.
@@ Appendix A Proofs
$\arg\max_{h_{1}}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{h_{1}}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}^{2}+\frac{1}{4}h^{2}_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]$
$=\arg\min_{h_{1}}E\left[\left(\frac{1}{2}h_{1}(X)-\left\{(1-A)\pi_{1}(X)-1\right\}\right)^{2}\right]$
$=2E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg].$
@@
$\breve{\pi}_{1}$ $=\arg\min_{\pi_{1}}\max_{h_{1}}E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]$
$=\arg\min_{\pi_{1}}E\left[2\left\{(1-A)\pi_{1}(X)-1\right\}E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg]-E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg]^{2}\right]$
$=\arg\min_{\pi_{1}}E\left[E\bigg[(1-A)\pi_{1}(X)-1\bigg|X\bigg]^{2}\right],$
@@
$E\bigg[(1-A)\breve{\pi}_{1}(X)-1\bigg|X\bigg]=0.$
@@
$E\left[\left\{(1-A)\breve{\pi}_{1}(X)-1\right\}h_{1}(X)\right]=E\left[E\bigg[(1-A)\breve{\pi}_{1}(X)-1\bigg|X\bigg]h_{1}(X)\right]=0.$
@@
$\arg\max_{h_{2}}E\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{h_{2}}E\left[\left\{A\pi_{2}(M,X)(1-A)\hat{\pi}_{1}(X)\right\}^{2}+\frac{1}{4}h^{2}_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)\right]$
$=\arg\min_{h_{2}}E\left[\left(\frac{1}{2}h_{2}(M,X)-\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\right)^{2}\right]$
$=2E\bigg[A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg|M,X\bigg].$
@@
$\breve{\pi}_{2}$ $=\arg\min_{\pi_{2}}\max_{h_{2}}E\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$=\arg\min_{\pi_{2}}E\bigg[2\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}E\big[A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\big|M,X\big]$
$~~~~~~~~~~~~~~~~~~~~-E\big[A\pi_{2}(M,X)-(1-A)\pi_{1}(X)\big|M,X\big]^{2}\bigg]$
$=\arg\min_{\pi_{2}}E\left[E\bigg[A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg|M,X\bigg]^{2}\right],$
@@
$E\bigg[A\breve{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\bigg|M,X\bigg]=0.$
@@
$E\left[\left\{A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\right\}h_{2}(M,X)\right]$
$=E\left[E\bigg[A\breve{\pi}_{2}(M,X)-(1-A)\pi_{1}(X)\bigg|M,X\bigg]h_{2}(M,X)\right]=0.$
@@
$\max_{h_{1}\in\mathcal{H}}E_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$=\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
@@
$E_{n}[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)]$ $=\sum_{i=1}^{n}h_{1}(X_{i})\{\xi_{n}(\pi_{1})\}_{i}={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\},$
$E_{n}[h_{1}^{2}(X)]$ $=\frac{1}{n}{\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}^{2}{\alpha^{\ast}},$
@@
$E_{n}\Big[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$={\alpha^{\ast}}^{\top}K_{\mathcal{H}_{1},n}\{\xi_{n}(\pi_{1})\}-{\alpha^{\ast}}^{\top}(\frac{1}{4n}K_{\mathcal{H}_{1},n}^{2}+\lambda_{\mathcal{H}_{1}}K_{\mathcal{H}_{1},n}){\alpha^{\ast}}.$
@@
$\alpha^{\ast}=\frac{1}{2}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\},$
@@
$\max_{h_{1}\in\mathcal{H}_{1}}E_{n}\Big[\{(1-A)\pi_{1}(X)-1\}h_{1}(X)-\frac{1}{4}h_{1}^{2}(X)\Big]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}$
$=\frac{1}{4}\{\xi_{n}(\pi_{1})\}^{\top}K_{\mathcal{H}_{1},n}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}\{\xi_{n}(\pi_{1})\}.$
@@
where $\Gamma_{1}=\frac{1}{4}K_{\mathcal{H}_{1},n}(\frac{1}{4n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$
@@
$=\min_{\alpha\in\mathbb{R}^{n}}\frac{1}{n^{2}}\alpha^{\top}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha-\frac{2}{n^{2}}e_{n}^{\top}\Gamma_{1}diag(1-A)K_{\Pi_{1},n}\alpha+\lambda_{\Pi_{1}}\alpha^{\top}K_{\Pi_{1},n}\alpha+c,$
@@
$\alpha=\Big(K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big)^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n}.$
@@
    $E\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right]=E\left[\left\{E[1-A\mid X]\pi_{1}(X)-1\right\}h_{1}(X)\right]=0.$